# Controllable Video Generation: A Survey

### v1

@@ 4.2.1 Person-Guided Generation
The primary input for person-guided control includes one or more ID images of the target person. These can be facial portraits or full Face and body synthesis. A driving signal—such as a text prompt, pose sequence, or audio—is also provided to guide motion. The ID images capture static visual identity, while the driving signal controls dynamics. Some methods, like ID-Animator[136], use a single image, while others, like Movie Weaver[131], rely on a set of images covering both face and body.
Representative diffusion-based methods explore various architectures and identity-injection strategies. EchoVideo[132] adopts a two-stage U-Net pipeline to decouple identity and motion. ID-Animator[136] integrates a ViT-based face adapter via cross-attention, while Movie Weaver[131] leverages a DiT backbone with anchor-concept prompts for tuning-free generation. SkyReels-A2[128] encodes multi-view ID and prompt concepts into a temporally-aware DiT. AnyCharV[130] enhances layout controllability through segmentation and bounding box guidance. PersonalVideo[129] applies a reward-based objective to a frozen LDM for identity-motion alignment. ConsisID[133] and Magic-Me[137] introduce frequency-aware and dynamic attention mechanisms, respectively, to reduce identity drift. Vlogger [138] combines audio-driven dynamics and identity conditioning using a dual-stream temporal transformer.
Nevertheless, key challenges remain. Sparse references limit generalization under novel views or occlusions, often leading to identity drift. Identity-motion balancing remains difficult, as stronger identity control may restrict expressiveness, while dominant motion cues can distort identity. Temporal consistency is also limited due to the frame-wise nature of diffusion models. Finally, real-world usability demands methods that are fast, tuning-free, and robust across diverse identities—goals not yet fully achieved.
@@ 4.2.2 Subject-Guided Generation
Subject-guided generation refers to the process of generating video content where the primary focus is on guiding the generation with specific subjects, such as particular characters, objects, or scenes, to ensure that the generated video adheres to user-defined subject constraints. This method enables the customization of content by explicitly controlling which subjects appear, how they interact, and in what context, offering more precise control over the generated content. Initial research often concentrated on single-subject customization, as exemplified by VideoBooth [376], which utilizes image prompts for enhanced content control, and DreamVideo [377], which decouples subject learning from motion learning. However, the field has rapidly progressed to address the more complex challenge of multi-subject video generation. Studies such as VideoDreamer [378], DisenStudio [157], CustomVideo [379], ConceptMaster [368], MovieWeaver [131], and VideoAlchemy [150] are at the forefront of addressing key difficulties in multi-subject scenarios. These challenges include ensuring the correct co-occurrence and temporally consistent appearance of multiple subjects, preserving their visual identities, avoiding attribute binding issues, and accurately assigning actions. Such approaches typically extend pretrained video diffusion models by incorporating several key innovations: novel fine-tuning strategies (e.g., VideoDreamer’s [378] Disen-Mix Finetuning, DisenStudio’s [157] motion-preserved disentangled fine-tuning); advancements in attention mechanisms (e.g., CustomVideo’s [379] attention control with object segmentation); and the development of specialized datasets (e.g., VideoDreamer’s [378] MultiStudioBench, CustomVideo’s [379] multi-subject dataset, and VideoAlchemy’s [150] and ConceptMaster’s [368] data construction pipelines).
@@
To further enhance the quality, controllability, and usability of customized videos, researchers are exploring various innovative approaches. In pursuit of model efficiency and generalization, zero-shot or tuning-free methods have emerged, exemplified by SUGAR [380], which leverages large-scale synthetic datasets, and VideoMaker [381], which utilizes the intrinsic feature extraction and injection capabilities of Video Diffusion Models (VDMs). ConsisID [382] offers a tuning-free method for identity preservation based on frequency decomposition. Still-Moving [383] uniquely facilitates the customization of video models using solely image data through the training of spatial adapters. Another key research area is fine-grained control over motion and concepts. For instance, MotionBooth [152] focuses on precise object and camera movement control. CustomTTT [147] employs test-time training to effectively combine multiple individually trained concepts (e.g., appearance and motion LoRAs), while CustomCrafter [384] aims to preserve the model’s intrinsic motion generation and concept composition capabilities during subject learning.
@@ 4.3 Image Control
The image-guided video generation aims to generate a video from a given reference image. This form of control is crucial for producing videos that are not only visually consistent with the reference but also maintain coherence across temporal frames. By conditioning the generation process on reference image features, models can generate video content with enhanced visual alignment, stylistic fidelity, and semantic relevance to the source image.
For some earlier works, Make It Move [207] and ImaGINator [208] propose novel Generative Adversarial Networks (GANs) architectures to generate video sequences from a reference image, while VideoGPT [209] introduces a VQ-VAE-based framework that surpasses the SOTA GANs models for the video generation task. To achieve a precise object motion expression, PiLife [204], LaMD [203], and AnimateAnything [385] optimize their framework from the mask, inversion, noise strategy, and the encoder structure, keeping the coherent and realistic motion of the generated video. Inspired by the SDXL [386], I2VGen-XL [200] proposes a two-stage training strategy, using static images as the primary condition to generate high-quality videos. Some works aimed to use the high-quality datasets to train a robust video generation framework. For instance, VideoCrafter1 [3] is trained on the LAION COCO 600M and Webvid10M dataset, and SVD [47] applies their proposed training scheme on a large video dataset that roughly contains 600 million samples. To alleviate the non-zero terminal signal-to-noise ratio, I4VGEN [179] incorporates image information into the inference process that reduces the costs and parameters significantly. Approaches such as TRIP [192], AtomoVideo [194], Tuning-free I2V [195], UNIVG [197], OmniTokenizer [189], Emu Video [184], FrameBridge [178] and I2V-Adapter [186] focus on addressing the insufficient consistency between the given image and the generated video, constructing a unified, high-quality video generation framework. To further construct a comprehensive video generation model, STIV [174] and Hunyuan Video [17] are based on the Diffusion Transformer (DiT) framework with the image condition, which perform better in both T2V and I2V tasks.
@@
Another challenge is to keep the alignment between the reference and the given prompt. Specifically, at the start of the generated video, the image always plays a more crucial role for the details, style, and the object location; however, in the later stage, the effects of the reference image can easily be weakened by the prompt. To solve this problem, DreamVideo [183] introduces an Image Retention module, maintaining both information from the input image and prompt. AID [190] incorporates an MLLM (Multimodal Large Language Model) and DQFormer (Dual Query Transformer) to predict future frames based on the given key frame. DynamiCrafter [180] utilizes a dual-stream image injection paradigm to adapt various applications (e.g., animation, storytelling). Techniques like Cond image leak [177] and EDG [170] aim to combat the issue of limited motion degrees and the unexpected motion with the prompt in the generated video. However, for the multi-object scenarios and the long-term video generation, the accuracy and consistency of the motion are also a challenge. Research like Through-The-Mask [173], Lumiere [176], TI2V-Zero [191], and EasyAnimate [187] attempt to utilize the powerful basic model (e.g., DiT) or improve the computation paradigm (e.g., compute all frames once and employ a special inversion strategy) while achieving a superior inference efficiency.
@@
Despite those remarkable achievements, image condition can also be easily combined with various conditions such as BBox, motion, and depth. Motion-I2V [196] first deduces the potential motion from the reference image, and then uses both predicted motion and image to generate high-quality video. Decoupled I2V [198] disentangles the motion vector and the image to obtain a memory-efficient and consistent video generation method. FrameGuidance [89] introduces a training-free strategy that supports flexible control (e.g., image, sketch, style). Similarly, MotionStone [175] also models the motion by the motion head and injects it into the Diffusion Transformer. Besides, MoVideo [185] utilizes the motion and keyframe with the depth and optical flow for video generation. ST-I2V [201] cropped several areas from the reference image, using them to maintain the semantic information of the video. DanceTogether [79], the first end-to-end framework for controllable multi-person video generation, consists of a MultiFace Encoder and MaskPoseAdapter, which enables precise identity-to-action alignment.
@@
Differently, PhysGen [181] consideres the physical properties like mass or elasticity, and external factors such as forces and environmental conditions. By giving input force and images as conditions, it can generate a video without a training process. HoloTime [205] introduces a 360-degree 4D scene generation framework to reconstruct high-quality 4D scene video. ConsistI2V [188] concatenates the input image to the noise and optimizes the attention calculation, not only performing better in consistency, but also supports various condition inputs (e.g., layout, camera). To achieve more user-friendly video generation, video doodles [202] and Follow-Your-Click [193] allow users to insert hand-drawn animations and provide a click to select which area to move, benefiting from better interactivity. Additionally, some work has explored using an image as the reference for other model architectures or tasks. NOVA [50, 387] reformulates the video generation task as a non-quantized autoregressive modeling of temporal and spatial prediction, achieving a superior performance with fewer parameters. VideoPanda [206] also introduces a long video generation framework using auto-regression. Structure and Content-Guided Video Synthesis [182] and Large-Motion Frame Interpolation [171] extend the image condition to the video editing and video frame interpolation applications, respectively.
@@ 4.4.1 Flow-Guided Generation
Flow guidance can be specified through optical flow fields, motion trajectories, or displacement representations, either explicitly computed or implicitly learned. For example, Motion-I2V [196] improves visual consistency by conditioning on motion trajectories, while MOFA-Video [211] animates static images using user-defined motion fields.
Recent approaches integrate flow-based guidance into generative architectures. I2VControl [210] ensures temporal coherence through disentangled flow representations, while MCDiff [212] employs motion-conditioned diffusion models for precise motion control. MOFA-Video [211] aligns motion fields for controllable animation, and Motion-I2V [196] emphasizes flow-based trajectory conditioning to enhance consistency.
Challenges include maintaining motion consistency in complex scenes, avoiding artifacts from inaccurate flow estimation, and efficiently integrating flow information without computational overhead. Techniques like MCDiff [212] and Motion-I2V [196] address these issues but highlight the need for further advancements in flow-guided video generation.
@@ 4.4.2 Trajectory-Guided Generation
Trajectory input can take various forms: a sequence of 2D or 3D coordinates $\{\mathbf{x}_{t},\mathbf{y}_{t},(\mathbf{z}_{t})\}_{t=1}^{T}$ for a point of interest, bounding box tracks $\{\text{bbox}_{t}\}_{t=1}^{T}$ defining the extent and location of an object over time (as in Boximator [125]), user-sketched paths in an initial frame, or even textual descriptions of motion (e.g., “move object A from left to right”). Drag-based interfaces, as explored in DragNuwa [239], DragAnything [238], DragEntity [234], C-Drag [227], and OmniDrag [231], allow interactive specification of start and end points for object parts.
Many recent methods leverage diffusion models for trajectory control. MotionBooth [152] and Motion-I2V [196] focus on generating video from an image based on motion prompts or trajectories. TrailBlazer [222] and Direct-a-Video [220] (which also handles camera control) enable explicit path following. Peekaboo [221] controls object appearance and disappearance along trajectories. Some methods, such as InTraGen [233] and MOFA-Video [211], focus on fine-grained trajectory control. Trajectory information is often encoded as a sequence of coordinates or heatmaps and is used to guide the cross-attention layers or directly modify the latent representations in diffusion models. Warping fields or flow-based methods can also be guided by trajectories. MCDiff [240] introduced motion-controlled diffusion. More advanced concepts like Perception-as-Control [226] and Motion Canvas [213] provide sophisticated frameworks. LeviTor [225] and LayerAnimate [388] explore layered representations for controllable animation along trajectories.
Key challenges include generating motion that is not only accurate to the specified trajectory but also appears natural and physically plausible, avoiding jerky or sliding movements. Ensuring that the guided object interacts realistically with other scene elements and the background is difficult. Handling complex, intersecting, or occluded trajectories for multiple objects (as in Collaborative Video Diffusion [219]) requires sophisticated reasoning. Maintaining long-range temporal consistency of the objectś appearance and the trajectory adherence over many frames is also crucial. FreeTraj [218], Tora [215], and MagicMotion [214] aim to address some of these complexities, with works like I2VControl [210] and ObjCtrl-2.5D [232] focusing on image-to-video with trajectory control.

### v2

@@ 5.2.1 Person-Guided Generation
The primary input for person-guided control includes one or more ID images of the target person. These can be facial portraits or full Face and body synthesis. A driving signal—such as a text prompt, pose sequence, or audio—is also provided to guide motion. The ID images capture static visual identity, while the driving signal controls dynamics. Some methods, like ID-Animator[he2024idanimator], use a single image, while others, like Movie Weaver[liang2025movieweaver] rely on a set of images covering both face and body.
Representative diffusion-based methods explore various architectures and identity-injection strategies. ID-Animator[he2024idanimator] integrates a ViT-based face adapter via cross-attention, while Movie Weaver[liang2025movieweaver] leverages a DiT backbone with anchor-concept prompts for tuning-free generation. PersonalVideo[li2025personalvideo] and Vlogger [zhuang2024vlogger] applie a reward-based objective to a frozen LDM for identity-motion alignment. ConsisID [yuan2024identitypreserving] and Magic-Me [ma2024magicme] introduce frequency-aware and dynamic attention mechanisms.
@@ 5.2.2 Subject-Guided Generation
Subject-guided generation [zhang2024ssr, jiang2023videobooth, song2025layertracer, song2025makeanything] refers to the process of generating video content where the primary focus is on guiding the generation with specific subjects, such as particular characters, objects, or scenes, to ensure that the generated video adheres to user-defined subject constraints. This method enables the customization of content by explicitly controlling which subjects appear, how they interact, and in what context, offering more precise control over the generated content. Initial research often concentrated on single-subject customization, as exemplified by VideoBooth [jiang2023videobooth], which utilizes image prompts for enhanced content control, and DreamVideo [DreamVideo], which decouples subject learning from motion learning. However, the field has rapidly progressed to address the more complex challenge of multi-subject video generation. Studies such as DisenStudio [chen2024disenstudio], CustomVideo [CustomVideo], and MovieWeaver [liang2025movieweaver]are at the forefront of addressing key difficulties in multi-subject scenarios. These challenges include ensuring the correct co-occurrence and temporally consistent appearance of multiple subjects, preserving their visual identities, avoiding attribute binding issues, and accurately assigning actions. Such approaches typically extend pretrained video diffusion models by incorporating several key innovations: novel fine-tuning strategies (e.g., VideoDreamer’s [chen2023videodreamer] Disen-Mix Finetuning, DisenStudio’s [chen2024disenstudio] motion-preserved disentangled fine-tuning); advancements in attention mechanisms (e.g., CustomVideo’s [CustomVideo] attention control with object segmentation); and the development of specialized datasets.
@@
To further enhance the quality, controllability, and usability of customized videos, researchers are exploring various innovative approaches. In pursuit of model efficiency and generalization, zero-shot or tuning-free methods have emerged, exemplified by SUGAR [SUGAR], which leverages large-scale synthetic datasets. ConsisID [yuan2024identitypreserving] offers a tuning-free method for identity preservation based on frequency decomposition. Still-Moving [Still-Moving] uniquely facilitates the customization of video models using solely image data through the training of spatial adapters. Another key research area is fine-grained control over motion and concepts. For instance, MotionBooth [wu2024motionbooth] focuses on precise object and camera movement control. CustomTTT [CustomTTT] employs test-time training to effectively combine multiple individually trained concepts (e.g., appearance and motion LoRAs), while CustomCrafter [CustomCrafter] aims to preserve the model’s intrinsic motion generation and concept composition capabilities during subject learning.
5.2.3 Summary
In this area, ID serves as a pivotal modality, utilizing conditional encoders and general control models (e.g., Adapters) to achieve the core goal of personalized customization. While UNet remains the dominant backbone (e.g., PersonalVideo [li2025personalvideo]), there is a strategic shift toward DiT (e.g., Movie Weaver [liang2025movieweaver]) to leverage global attention for improved temporal consistency.
Driven by the user demand for efficiency, the technical roadmap is transitioning from slow fine-tuning to Tuning-free approaches (e.g., ConsisID [yuan2024identitypreserving]), which directly aligns with the users’ demand for real-time. However, current limitations in multi-subject generation reveal a gap in the unified system of creative supports, underscoring the future need for advanced disentanglement and composition to resolve conflicts between different controllable methods (e.g., conditional encoders and motion modules) without performance degradation.
@@ 5.3 Image Control
The image-guided video generation aims to generate a video from a given reference image. This form of control is crucial for producing videos that are not only visually consistent with the reference but also maintain coherence across temporal frames. By conditioning the generation process on reference image features, models can generate video content with enhanced visual alignment, stylistic fidelity, and semantic relevance to the source image. Approaches such as TRIP [Zhang_2024_CVPR], OmniTokenizer [wang2024omnitokenizerjointimagevideotokenizer], Emu Video [girdhar2024emuvideofactorizingtexttovideo], and I2V-Adapter [10.1145/3641519.3657407] focus on addressing the insufficient consistency between the given image and the generated video, constructing a unified, high-quality video generation framework. To further construct a comprehensive video generation model, STIV [lin2024stivscalabletextimage] and Hunyuan Video [kong2025hunyuanvideosystematicframeworklarge] are based on the Diffusion Transformer (DiT) framework with the image condition, which perform better in both T2V and I2V tasks.
@@
Another challenge is to keep the alignment between the reference and the given prompt. Specifically, at the start of the generated video, the image always plays a more crucial role for the details, style, and the object location; however, in the later stage, the effects of the reference image can easily be weakened by the prompt. To solve this problem, DreamVideo [10887583] introduces an Image Retention module, maintaining both information from the input image and prompt. AID [xing2024aidadaptingimage2videodiffusion] incorporates an MLLM (Multimodal Large Language Model) and DQFormer (Dual Query Transformer) to predict future frames based on the given key frame. DynamiCrafter [10.1007/978-3-031-72952-2_23] utilizes a dual-stream image injection paradigm to adapt various applications (e.g., animation, storytelling). Techniques like Cond image leak [zhao2024identifyingsolvingconditionalimage] and EDG [tian2025extrapolatingdecouplingimagetovideogeneration] aim to combat the issue of limited motion degrees and the unexpected motion with the prompt in the generated video. However, for the multi-object scenarios and the long-term video generation, the accuracy and consistency of the motion are also a challenge. Research like Through-The-Mask [yariv2025throughthemaskmaskbasedmotiontrajectories], Lumiere [10.1145/3680528.3687614], TI2V-Zero [ni2024ti2vzerozeroshotimageconditioning], and EasyAnimate [xu2024easyanimatehighperformancelongvideo] attempt to utilize the powerful basic model (e.g., DiT) or improve the computation paradigm (e.g., compute all frames once and employ a special inversion strategy) while achieving a superior inference efficiency.
@@
Despite those remarkable achievements, image condition can also be easily combined with various conditions such as BBox, motion, and depth. Motion-I2V [10.1145/3641519.3657497] first deduces the potential motion from the reference image, and then uses both predicted motion and image to generate high-quality video. Decoupled I2V [shen2023decouplecontentmotionconditional] disentangles the motion vector and the image to obtain a memory-efficient and consistent video generation method. Similarly, MotionStone [shi2024motionstonedecoupledmotionintensity] also models the motion by the motion head and injects it into the Diffusion Transformer. Besides, MoVideo [liang2024movideomotionawarevideogeneration] utilizes the motion and keyframe with the depth and optical flow for video generation. ST-I2V [10.1145/3581783.3611897] cropped several areas from the reference image, using them to maintain the semantic information of the video.
@@
Differently, PhysGen [liu2024physgenrigidbodyphysicsgroundedimagetovideo] consideres the physical properties like mass or elasticity, and external factors such as forces and environmental conditions. By giving input force and images as conditions, it can generate a video without a training process. HoloTime [zhou2025holotimetamingvideodiffusion] introduces a 360-degree 4D scene generation framework to reconstruct high-quality 4D scene video. ConsistI2V [ren2024consisti2venhancingvisualconsistency] concatenates the input image to the noise and optimizes the attention calculation, not only performing better in consistency, but also supports various condition inputs (e.g., layout, camera). To achieve more user-friendly video generation, video doodles [10.1145/3592413] and Follow-Your-Click [ma2024followyourclickopendomainregionalimage] allow users to insert hand-drawn animations and provide a click to select which area to move, benefiting from better interactivity. Additionally, some work has explored using an image as the reference for other model architectures or tasks. NOVA [deng2025autoregressivevideogenerationvector] reformulates the video generation task as a non-quantized autoregressive modeling of temporal and spatial prediction, achieving a superior performance with fewer parameters. VideoPanda [xie2025videopandavideopanoramicdiffusion] also introduces a long video generation framework using auto-regression. Structure and Content-Guided Video Synthesis [Esser_2023_ICCV] extends the image condition to the video editing and video frame interpolation applications, respectively.
In summary, image control utilizes conditional encoders to map visual features into the backbone, fulfilling the user need for personalization and customization through high-fidelity visual alignment. In terms of the foundation models, the framework highlights a dynamic architectural ecosystem. Traditional UNet structures (e.g., I2V-Adapter [10.1145/3641519.3657407]) are increasingly being complemented by DiT (e.g., STIV [lin2024stivscalabletextimage]) and Autoregressive models (e.g., NOVA [deng2025autoregressivevideogenerationvector]), enabling superior scalability and temporal coherence for complex generation tasks. However, persistent challenges in multi-object scenarios and prompt-image alignment suggest a continued need for better explainability to help users understand how models balance conflicting inputs during the generation process.
@@ 5.4.1 Flow-Guided Generation
Flow guidance can be specified through optical flow fields, motion trajectories, or displacement representations, either explicitly computed or implicitly learned. For example, Motion-I2V [10.1145/3641519.3657497] improves visual consistency by conditioning on motion trajectories, while MOFA-Video [niu2024mofavideo] animates static images using user-defined motion fields.
Recent approaches integrate flow-based guidance into generative architectures. I2VControl [feng2024i2vcontrol] ensures temporal coherence through disentangled flow representations, while MCDiff [chen2023motionconditioneddiffusionmodelcontrollable] employs motion-conditioned diffusion models for precise motion control. MOFA-Video [niu2024mofavideo] aligns motion fields for controllable animation, and Motion-I2V [10.1145/3641519.3657497] emphasizes flow-based trajectory conditioning to enhance consistency.
Challenges include maintaining motion consistency in complex scenes, avoiding artifacts from inaccurate flow, and efficiently integrating flow information without computational overhead. Techniques like MCDiff [chen2023motionconditioneddiffusionmodelcontrollable] and Motion-I2V [10.1145/3641519.3657497] address these issues but highlight the need for further advancements in flow-guided video generation.
Table 5: Quantitative comparison of camera-guided video generation on 3D static and 4D dynamic scenes, using public and internet data. We evaluate generation quality (FID, $\mathrm{CLIP}_{\mathrm{sim}}$ for static; FVD, $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}$ for dynamic) and trajectory accuracy (ATE, RPE-T, RPE-R). All methods use official code with identical inputs. $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better. The best and secondary results are in bold and underline, respectively.
Methods Generation Quality Trajectory Accuracy
Static Dynamic Static Dynamic
FID $\downarrow$ $\mathrm{CLIP}_{\mathrm{sim}}\uparrow$ FVD $\downarrow$ $\mathrm{CLIP\text{-}V}_{\mathrm{sim}}\uparrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$ ATE $\downarrow$ RPE-T $\downarrow$ RPE-R $\downarrow$
See3D [ma2025you] 123.26 0.941 0.091 0.089 0.250
ViewCrafter [yu2024viewcrafter] 117.50 0.930 0.236 0.315 0.728
ViewExtrapolator [liu2024novel] 125.50 0.930 108.48 0.913 0.183 0.260 0.882 1.040 1.208 4.750
TrajectoryAttention [xiao2024trajectoryattention] 122.37 0.920 106.94 0.911 0.159 0.238 0.532 0.605 1.238 3.560
TrajectoryCrafter [yu2025trajectorycrafter] 111.49 0.910 97.31 0.923 0.090 0.152 0.267 0.431 1.078 8.950
NVS-Solver [you2024nvs] 118.64 0.937 0.224 0.268 1.056
WorldForge[song2025worldforge] 96.08 0.948 93.17 0.938 0.077 0.086 0.221 0.527 0.826 2.690
@@ 5.4.2 Trajectory-Guided Generation
Trajectory input can take various forms: a sequence of 2D or 3D coordinates $\{\mathbf{x}_{t},\mathbf{y}_{t},(\mathbf{z}_{t})\}_{t=1}^{T}$ for a point of interest, bounding box tracks $\{\text{bbox}_{t}\}_{t=1}^{T}$ defining the extent and location of an object over time (as in Boximator [wang2024boximator]), user-sketched paths in an initial frame, or even textual descriptions of motion (e.g., “move object A from left to right”). Drag-based interfaces, as explored in DragAnything [wu2024draganything] and DragEntity [wan2024dragentity], allow interactive specification of start and end points for object parts.
Many recent methods leverage diffusion models for trajectory control. MotionBooth [wu2024motionbooth] and Motion-I2V [10.1145/3641519.3657497] focus on generating video from an image based on motion prompts or trajectories. TrailBlazer [ma2024trailblazer] and Direct-a-Video [yang2024directavideo] (which also handles camera control) enable explicit path following. Peekaboo [jain2024peekaboo] controls object appearance and disappearance along trajectories. MOFA-Video [niu2024mofavideo] focuses on fine-grained trajectory control. Trajectory information is often encoded as a sequence of coordinates or heatmaps and is used to guide the cross-attention layers or directly modify the latent representations in diffusion models. Warping fields or flow-based methods can also be guided by trajectories. MCDiff [chen2023motionconditioneddiffusionmodelcontrollable] introduced motion-controlled diffusion. More advanced concepts like Perception-as-Control [chen2025perceptionascontrol] and Motion Canvas [xing2025motioncanvas] provide sophisticated frameworks. LeviTor [wang2025levitor] and LayerAnimate [yang2025layeranimate] explore layered representations for controllable animation along trajectories.
Key challenges include generating motion that is not only accurate to the specified trajectory but also appears natural and physically plausible, avoiding jerky or sliding movements. Ensuring that the guided object interacts realistically with other scene elements and the background is difficult. Handling complex, intersecting, or occluded trajectories for multiple objects (as in Collaborative Video Diffusion [kuang2024collaborative]) requires sophisticated reasoning. Maintaining long-range temporal consistency of the objectś appearance and the trajectory adherence over many frames is also crucial. Tora [zhang2025tora] and MagicMotion [li2025magicmotion] aim to address some of these complexities, with works like I2VControl [feng2024i2vcontrol] focusing on image-to-video with trajectory control.